On February 1, 2000, Herro Corp. sold an $800 million bond issue to finance the purchase of a new distribution facility. These bonds were issued in $1,000 denominations with a maturity date of February 1, 2030. The bonds have a coupon rate of 10.00% with interest paid semiannually.

Required: a) Determine the value today, February 1, 2020 of one of these bonds to an investor who requires an 8 percent return on these bonds. Why is the value today different from the par value?

b) Assume that the bonds are selling for $1,175.00. Determine the current yield and the yield-to-maturity.
Answer #1

(a) Price of one bond today= $1,135.90

Details as follows:

po E14 f =PV(E13, E10, E11,E3)*-1 в ср FV Cell Ref. Values $1,000 10% Semiannual 1 2 Formula 3 Face Value (FV) Given 4 Coupon

Price is higher than the par value because the rate of return is lower than the coupon rate. Rate of return (YTM) is the discount rate which the cash flows are discounted in order to arrive at the present value. Cash flows, on the other hand, comprise of coupon payments at the originally fixed higher rate and the face value on redemption which also remains unchanged.

(b) If the bonds are selling at $1,175:

Current yield= (Interest per year/Price)*100= ($100/$1,175)*100 = 8.5106383%

Yield to Maturity (YTM) = 7.483218%  Details as follows:

А D Clipboard Font Alignment D12 f =RATE(D9,010,D11*-1,D2)*D5 Bc 1 Formula 2 Face Value (FV) Given FV $1,000 3 Coupon rate (R